Janine is 5 years older than three times her sister’s age. Together, their ages sum to 17. How old is her sister?
Answer by ikleyn(52799)

.

x = sister;

then Janine = 3x + 5.


Together


    x + (3x+5) = 17.

    4x = 17 - 5 = 12

     x          = 12/4 = 3.


ANSWER.  The sister is 3 years old.  Janine is 17-3 = 14 years old.

Solved.